OpenStudy (anonymous):

Would anyone care to show me how to draw parabolas :s? I have a link to a whiteboard sharing site. I'd be much appreciated, thanks c:

jimthompson5910 (jim_thompson5910):

Do you have an equation to work with? If so, then simply plug in values of x to get corresponding values of y. From there, you plot the points. Afterwards, you draw a curve through the points to graph the parabola. The more points you plot, the more accurate your drawing is.

jimthompson5910 (jim_thompson5910):

Ex: Graphing y = x^2 When x = -2, y is y = 4 giving you (-2,4) When x = -1, y is y = 1 giving you (-1,1) etc You then plot these points and draw a curve through them

OpenStudy (anonymous):

generic parabola that faces up with vertex at the origin (0,0) is given by \(y=x^2\) http://www.wolframalpha.com/input/?i=y%3Dx^2

OpenStudy (anonymous):

you can shift this one up, down, left and right. if you are studying conic section you might have been given a formula \((y-k)^2=4p(x-h)\) or \[4p(y-k)=(x-h)^2\]